OpenStudy (anonymous):

If x + 1/x = 5, find the value of x^5 + 1/x^5

OpenStudy (anonymous):

well first find the value of x with the first equation then plug it in and ur good to go..

OpenStudy (anonymous):

\[\LARGE\frac{x + 1}{x} = 5 \implies 5x = x + 1 \implies 4x = 1 \implies x = \frac14\]Now, plug it into the second equation. \[\LARGE\frac{x^5 + 1}{x^5} \implies \frac{(\frac14)^5 + 1}{(\frac14)^5} \implies ?\]

OpenStudy (anonymous):

@Calcmathlete it is x+(1/x) isnt it???

OpenStudy (anonymous):

Is it?

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

Not sure...the question was a bit vague...

OpenStudy (anonymous):

i thought Calcmathlete was right.. its how i did it

OpenStudy (anonymous):

 That is what I meant. The question wants me to find the value without finding the value of x...

OpenStudy (anonymous):

welll then the prob is just raised to the 5th so whats 5 times its self 5 times

OpenStudy (anonymous):

@katiebugg That wouldn't work because if you raise it to the fifth power, it becomes: \[(x + \frac1x)^5 = 5^5\]

OpenStudy (anonymous):

but isnt every thing raised to the fifth except 1 that is but that wouldnt matter cuz its still 1